# Improving robustness to corruptions with multiplicative weight perturbations

### v1

@@ 2 Data Augmentation via Multiplicative Perturbations
$\min_{\boldsymbol{\omega}}\max_{||\boldsymbol{\xi}||_{2}\leq\rho}\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\xi};\mathcal{S})+\frac{\lambda}{2}||\boldsymbol{\omega}||_{F}^{2}$ (17)
which can be interpreted as optimizing DNNs under adversarial additive perturbations. To efficiently solve this problem, Foret et al. (2021) devise a two-step procedure for each iteration $t$:
$\boldsymbol{\xi}^{(t)}=\rho\frac{\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega}^{(t)};\mathcal{S})}{\left|\left|\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega}^{(t)};\mathcal{S})\right|\right|_{2}},\qquad\boldsymbol{\omega}^{(t+1)}=\boldsymbol{\omega}^{(t)}-\eta_{t}\left(\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega}^{(t)}+\boldsymbol{\xi}^{(t)};\mathcal{S})+\lambda\boldsymbol{\omega}^{(t)}\right)$ (18)
@@
$\min_{\boldsymbol{\omega}}\max_{||T_{\boldsymbol{\omega}}^{-1}\boldsymbol{\xi}||_{2}\leq\rho}\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\xi};\mathcal{S})+\frac{\lambda}{2}||\boldsymbol{\omega}||_{F}^{2}$ (19)
@@
$\hat{\boldsymbol{\xi}}^{(t)}=\rho\frac{\left(\boldsymbol{\omega}^{(t)}\right)^{2}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega}^{(t)};\mathcal{S})}{\left|\left|\boldsymbol{\omega}^{(t)}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega}^{(t)};\mathcal{S})\right|\right|_{2}},\quad\boldsymbol{\omega}^{(t+1)}=\boldsymbol{\omega}^{(t)}-\eta_{t}\left(\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega}^{(t)}+\hat{\boldsymbol{\xi}}^{(t)};\mathcal{S})+\lambda\boldsymbol{\omega}^{(t)}\right)$ (20)
@@
Algorithm 1 minimizes the expected loss in Eq. 16. Instead, we could minimize the loss under the adversarial multiplicative weight perturbation:
$\mathcal{L}_{\mathrm{max}}(\boldsymbol{\omega};\mathcal{S})\stackrel{{\scriptstyle\mathrm{def}}}{{=}}\max_{||\boldsymbol{\xi}||_{2}\leq\rho}\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\omega}\circ\boldsymbol{\xi};\mathcal{S})$ (21)
Similar to Foret et al. (2021), we solve this optimization problem by using a first-order Taylor expansion of $\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\omega}\circ\boldsymbol{\xi};\mathcal{S})$ to find an approximate solution of the inner maximization:
$\operatorname*{arg\,max}_{||\boldsymbol{\xi}||_{2}\leq\rho}\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\omega}\circ\boldsymbol{\xi};\mathcal{S})\approx\operatorname*{arg\,max}_{||\boldsymbol{\xi}||_{2}\leq\rho}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})+\left\langle\boldsymbol{\omega}\circ\boldsymbol{\xi},\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})\right\rangle$ (22)
@@
$\hat{\boldsymbol{\xi}}(\boldsymbol{\omega})=\rho\frac{\boldsymbol{\omega}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})}{||\boldsymbol{\omega}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})||_{2}}$ (23)
@@
$\nabla_{\boldsymbol{\omega}}\mathcal{L}_{\mathrm{max}}(\boldsymbol{\omega};\mathcal{S})$ $\approx\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\omega}\circ\hat{\boldsymbol{\xi}}(\boldsymbol{\omega});\mathcal{S})=\nabla_{\boldsymbol{\omega}}\left(\boldsymbol{\omega}+\boldsymbol{\omega}\circ\hat{\boldsymbol{\xi}}(\boldsymbol{\omega})\right)\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})\big{|}_{\boldsymbol{\omega}+\boldsymbol{\omega}\circ\hat{\boldsymbol{\xi}}(\boldsymbol{\omega})}$ (24)
$=\nabla_{\hat{\boldsymbol{\omega}}}\mathcal{L}(\hat{\boldsymbol{\omega}};\mathcal{S})+\nabla_{\boldsymbol{\omega}}\left(\boldsymbol{\omega}\circ\hat{\boldsymbol{\xi}}(\boldsymbol{\omega})\right)\nabla_{\hat{\boldsymbol{\omega}}}\mathcal{L}(\hat{\boldsymbol{\omega}};\mathcal{S})$ (25)
where $\hat{\boldsymbol{\omega}}$ is the perturbed weight:
$\hat{\boldsymbol{\omega}}=\boldsymbol{\omega}+\boldsymbol{\omega}\circ\hat{\boldsymbol{\xi}}(\boldsymbol{\omega})=\boldsymbol{\omega}+\frac{\boldsymbol{\omega}^{2}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})}{||\boldsymbol{\omega}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})||_{2}}$ (26)
By omitting the second summand in Eq. 25, we arrive at the gradient formula in the update rule of ASAM in Eq. 20. We have thus established a connection between ASAM and adversarial multiplicative weight perturbations.
@@ 4 Empirical evaluation
In this section, we assess the corruption robustness of DAMP and ASAM in image classification tasks. We conduct experiments using the CIFAR-10/100 (Krizhevsky, 2009), TinyImageNet (Le and Yang, 2015), and ImageNet (Deng et al., 2009) datasets. For evaluation on corrupted images, we utilize the CIFAR-10/100-C, TinyImageNet-C, and ImageNet-C datasets provided by Hendrycks and Dietterich (2019), as well as the ImageNet-$\bar{\text{C}}$ dataset provided by Mintun et al. (2021). The datasets from Hendrycks and Dietterich (2019) contain 15 common image corruptions, while ImageNet-$\bar{\text{C}}$ offers 10 additional corruptions. All corruptions in these datasets have 5 levels of severity. In terms of architectures, we use ResNet18 (He et al., 2016a) for CIFAR-10/100, PreActResNet18 (He et al., 2016b) for TinyImageNet, and ResNet50 (He et al., 2016a) and ViT-S/16 (Dosovitskiy et al., 2021) for ImageNet. We ran all experiments on a single machine with 8 Nvidia V100 GPUs. Detailed information for each experiment is provided in Appendix D.
@@ Results
We visualize the results for the ResNet18/CIFAR-100 setting in Fig. 1. The results for the ResNet18/CIFAR-10 and PreActResNet18/TinyImageNet settings are presented in Figs. 4 and 5 in the Appendix. Figs. 1, 4 and 5 demonstrate that DAMP improves predictive accuracy over plain SGD across all corruptions without compromising accuracy on clean images. Although Fig. 1 indicates that including zoom_blur as an augmentation when training ResNet18 on CIFAR-100 yields better results than DAMP on average, it also reduces accuracy on clean images and the brightness corruption. Overall, these figures show that incorporating a specific corruption as data augmentation during training enhances robustness to that particular corruption but may reduce performance on clean images and other corruptions. In contrast, DAMP consistently improves robustness across all corruptions. Notably, DAMP even enhances accuracy on clean images in the PreActResNet18/TinyImageNet setting, as shown in Fig. 5.
@@ 4.2 Comparing DAMP to random additive perturbations
In this section, we investigate whether additive weight perturbations can also enhance corruption robustness. To this end, we compare DAMP with its variant, Data Augmentation via Additive Perturbations (DAAP). Unlike DAMP, DAAP perturbs weights during training with random additive Gaussian noises centered at 00, as detailed in Algorithm 3 in the Appendix. Fig. 2 presents the results of DAMP and DAAP under different noise standard deviations, alongside standard SGD. Overall, Fig. 2 shows that across different experimental settings, the corruption robustness of DAAP is only slightly better than SGD and is worse than DAMP. Therefore, we conclude that multiplicative weight perturbations are better than their additive counterparts at improving robustness to corruptions.
@@ 4.3 Benchmark results
In this section, we compare DAMP with SGD, SAM (Foret et al., 2021), and ASAM (Kwon et al., 2021). For SAM and ASAM, we optimize the neighborhood size $\rho$ by using $10\%$ of the training set as a validation set. Similarly, we adjust the noise standard deviation $\sigma$ for DAMP following the same procedure. For hyperparameters and additional training details, please refer to Appendix D.
@@
Fig. 3 visualizes the predictive errors of DAMP and the baseline methods on CIFAR-10/100 and TinyImageNet, with all methods trained for the same number of epochs. It demonstrates that DAMP consistently outperforms SGD across various datasets and corruption severities, despite having the same training cost. Notably, DAMP outperforms SAM under most corruption scenarios, even though SAM takes twice as long to train and has higher accuracy on clean images. Additionally, DAMP improves accuracy on clean images over SGD on CIFAR-100 and TinyImageNet. Finally, ASAM consistently surpasses other methods on both clean and corrupted images, as it employs adversarial multiplicative weight perturbations (Section 3). However, like SAM, each ASAM experiment takes twice as long as DAMP given the same epoch counts.
@@
Table 1 presents the predictive errors for the ResNet50 / ImageNet setting. It shows that DAMP consistently outperforms SAM under all corruption scenarios, despite having half the training cost and lower accuracy on clean images. Notably, DAMP surpasses the baselines on corrupted images in most cases, except for severity levels 1, 2, and 3 of ImageNet-C.
Table 1: DAMP surpasses the baselines on corrupted images in most cases. We report the predictive errors (lower is better) averaged over 3 seeds for the ResNet50 / ImageNet experiments. For each level of severity in ImageNet-C and ImageNet-$\bar{\text{C}}$, we report the average error over all corruption types. We use 90 epochs and the basic Inception-style preprocessing for all experiments.
Method ImageNet ImageNet-C ImageNet-$\bar{\text{C}}$
1 2 3 4 5 1 2 3 4 5
SGD $23.6_{\pm 0.2}$ $40.2_{\pm 0.1}$ $51.0_{\pm 0.2}$ $61.1_{\pm 0.3}$ $73.5_{\pm 0.2}$ $82.9_{\pm<0.1}$ $45.7_{\pm 0.1}$ $54.7_{\pm 0.1}$ $61.8_{\pm<0.1}$ $70.3_{\pm 0.1}$ $75.7_{\pm 0.1}$
DAMP $23.8_{\pm<0.1}$ $38.2_{\pm<0.1}$ $48.1_{\pm<0.1}$ $57.4_{\pm 0.2}$ $\bf{69.5_{\pm 0.3}}$ $\bf{79.9_{\pm 0.2}}$ $\bf{42.4_{\pm 0.1}}$ $\bf{51.2_{\pm 0.1}}$ $\bf{58.6_{\pm<0.1}}$ $\bf{67.8_{\pm 0.1}}$ $\bf{73.4_{\pm 0.1}}$
SAM $23.1_{\pm 0.1}$ $38.7_{\pm 0.3}$ $49.1_{\pm 0.3}$ $59.1_{\pm 0.3}$ $71.6_{\pm 0.1}$ $81.7_{\pm 0.1}$ $44.5_{\pm<0.1}$ $53.5_{\pm 0.1}$ $60.8_{\pm<0.1}$ $69.4_{\pm<0.1}$ $75.1_{\pm 0.1}$
ASAM $\bf{22.8_{\pm 0.1}}$ $\bf{37.4_{\pm 0.1}}$ $\bf{47.5_{\pm<0.1}}$ $\bf{57.3_{\pm 0.1}}$ $70.0_{\pm 0.1}$ $80.7_{\pm 0.1}$ $42.5_{\pm 0.1}$ $51.4_{\pm 0.1}$ $59.0_{\pm 0.1}$ $68.2_{\pm 0.1}$ $74.0_{\pm<0.1}$
ViT-S16 / ImageNet

### v2

@@ 2.2 Multiplicative weight perturbations simulate input corruptions
$\min_{\boldsymbol{\omega}}\max_{||\boldsymbol{\xi}||_{2}\leq\rho}{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\xi};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{\lambda}{2}||\boldsymbol{\omega}||_{F}^{2}}$ (8)
which can be interpreted as optimizing DNNs under adversarial additive weight perturbations. To efficiently solve this problem, Foret et al. (2021) devise a two-step procedure for each iteration $t$:
$\boldsymbol{\xi}^{(t)}=\rho\frac{\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega}^{(t)};\mathcal{S})}{\left|\left|\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega}^{(t)};\mathcal{S})\right|\right|_{2}},\qquad\boldsymbol{\omega}^{(t+1)}=\boldsymbol{\omega}^{(t)}-\eta_{t}\left({\color[rgb]{0.99609375,0.26171875,0.234375}\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega}^{(t)}+\boldsymbol{\xi}^{(t)};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\lambda\boldsymbol{\omega}^{(t)}}\right)$ (9)
@@
$\min_{\boldsymbol{\omega}}\max_{||T_{\boldsymbol{\omega}}^{-1}\boldsymbol{\xi}||_{2}\leq\rho}{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\xi};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{\lambda}{2}||\boldsymbol{\omega}||_{F}^{2}}$ (10)
@@
$\widehat{\boldsymbol{\xi}}^{(t)}=\rho\frac{\left(\boldsymbol{\omega}^{(t)}\right)^{2}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega}^{(t)};\mathcal{S})}{\left|\left|\boldsymbol{\omega}^{(t)}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega}^{(t)};\mathcal{S})\right|\right|_{2}},\quad\boldsymbol{\omega}^{(t+1)}=\boldsymbol{\omega}^{(t)}-\eta_{t}\left({\color[rgb]{0.99609375,0.26171875,0.234375}\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega}^{(t)}+\widehat{\boldsymbol{\xi}}^{(t)};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\lambda\boldsymbol{\omega}^{(t)}}\right)$ (11)
@@
Algorithm 1 minimizes the expected loss in Eq. 7. Instead, we could minimize the loss under the adversarial MWP:
${\color[rgb]{0.63671875,0.140625,0.6796875}\mathcal{L}_{\mathrm{max}}(\boldsymbol{\omega};\mathcal{S})}\vcentcolon=\max_{||\boldsymbol{\xi}||_{2}\leq\rho}{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\omega}\circ\boldsymbol{\xi};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\frac{\lambda}{2}||\boldsymbol{\omega}||_{F}^{2}}$ (12)
Following Foret et al. (2021), we solve this optimization problem by using a first-order Taylor expansion of $\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\omega}\circ\boldsymbol{\xi};\mathcal{S})$ to find an approximate solution of the inner maximization:
$\operatorname*{arg\,max}_{||\boldsymbol{\xi}||_{2}\leq\rho}{\color[rgb]{0.99609375,0.26171875,0.234375}\mathcal{L}(\boldsymbol{\omega}+\boldsymbol{\omega}\circ\boldsymbol{\xi};\mathcal{S})}\approx\operatorname*{arg\,max}_{||\boldsymbol{\xi}||_{2}\leq\rho}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})+\left\langle\boldsymbol{\omega}\circ\boldsymbol{\xi},\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})\right\rangle$ (13)
@@
$\widehat{\boldsymbol{\xi}}(\boldsymbol{\omega})=\rho\frac{\boldsymbol{\omega}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})}{||\boldsymbol{\omega}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})||_{2}}$ (14)
@@
${\color[rgb]{0.63671875,0.140625,0.6796875}\nabla_{\boldsymbol{\omega}}\mathcal{L}_{\mathrm{max}}(\boldsymbol{\omega};\mathcal{S})}$ $\approx{\color[rgb]{0.99609375,0.26171875,0.234375}\nabla_{\boldsymbol{\omega}}\mathcal{L}(\widehat{\boldsymbol{\omega}};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\lambda\boldsymbol{\omega}}={\color[rgb]{0.99609375,0.26171875,0.234375}\nabla_{\boldsymbol{\omega}}\widehat{\boldsymbol{\omega}}\cdot\nabla_{\widehat{\boldsymbol{\omega}}}\mathcal{L}(\widehat{\boldsymbol{\omega}};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\lambda\boldsymbol{\omega}}$ (15)
$={\color[rgb]{0.99609375,0.26171875,0.234375}\nabla_{\widehat{\boldsymbol{\omega}}}\mathcal{L}(\widehat{\boldsymbol{\omega}};\mathcal{S})}+{\color[rgb]{1,0.5390625,0.55078125}\nabla_{\boldsymbol{\omega}}\left(\boldsymbol{\omega}\circ\widehat{\boldsymbol{\xi}}(\boldsymbol{\omega})\right)\cdot\nabla_{\widehat{\boldsymbol{\omega}}}\mathcal{L}(\widehat{\boldsymbol{\omega}};\mathcal{S})}+{\color[rgb]{0.03515625,0.5859375,0.9375}\lambda\boldsymbol{\omega}}$ (16)
where $\widehat{\boldsymbol{\omega}}$ is the perturbed weight:
$\widehat{\boldsymbol{\omega}}=\boldsymbol{\omega}+\boldsymbol{\omega}\circ\widehat{\boldsymbol{\xi}}(\boldsymbol{\omega})=\boldsymbol{\omega}+\rho\frac{\boldsymbol{\omega}^{2}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})}{||\boldsymbol{\omega}\circ\nabla_{\boldsymbol{\omega}}\mathcal{L}(\boldsymbol{\omega};\mathcal{S})||_{2}}$ (17)
Similar to Foret et al. (2021), we omit the second summand in Eq. 16 for efficiency, as it requires calculating the Hessian of the loss. We then arrive at the gradient formula in the update rule of ASAM in Eq. 11. We have thus established a connection between ASAM and adversarial MWPs.
@@ 4 Empirical evaluation
In this section, we assess the corruption robustness of DAMP and ASAM in image classification tasks. We conduct experiments using the CIFAR-10/100 (Krizhevsky, 2009), TinyImageNet (Le and Yang, 2015), and ImageNet (Deng et al., 2009) datasets. For evaluation on corrupted images, we utilize the CIFAR-10/100-C, TinyImageNet-C, and ImageNet-C datasets provided by Hendrycks and Dietterich (2019), as well as ImageNet-$\bar{\text{C}}$ (Mintun et al., 2021), ImageNet-D (Zhang et al., 2024), ImageNet-A (Hendrycks et al., 2021), ImageNet-Sketch (Wang et al., 2019), ImageNet-Drawing (Salvador and Oberman, 2022), and ImageNet-Cartoon (Salvador and Oberman, 2022) datasets, which encapsulate a wide range of corruptions. Detail descriptions of these datasets are provided in Appendix E. We further evaluate the models on adversarial examples generated by the Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2014). In terms of architectures, we use ResNet18 (He et al., 2016a) for CIFAR-10/100, PreActResNet18 (He et al., 2016b) for TinyImageNet, ResNet50 (He et al., 2016a), ViT-S/16, and ViT-B/16 (Dosovitskiy et al., 2021) for ImageNet. We ran all experiments on a single machine with 8 Nvidia V100 GPUs. Appendix F includes detailed information for each experiment.
@@ Results
We visualize the results for the ResNet18/CIFAR-100 setting in Fig. 3. The results for the ResNet18/CIFAR-10 and PreActResNet18/TinyImageNet settings are presented in Figs. 5 and 6 in the Appendix. Figs. 3, 5 and 6 demonstrate that DAMP improves predictive accuracy over plain SGD across all corruptions without compromising accuracy on clean images. Although Fig. 3 indicates that including zoom_blur as an augmentation when training ResNet18 on CIFAR-100 yields better results than DAMP on average, it also reduces accuracy on clean images and the brightness corruption. Overall, these figures show that incorporating a specific corruption as data augmentation during training enhances robustness to that particular corruption but may reduce performance on clean images and other corruptions. In contrast, DAMP consistently improves robustness across all corruptions. Notably, DAMP even enhances accuracy on clean images in the PreActResNet18/TinyImageNet setting, as shown in Fig. 6.
@@ 4.2 Comparing DAMP to random additive perturbations
In this section, we investigate whether additive weight perturbations can also enhance corruption robustness. To this end, we compare DAMP with its variant, Data Augmentation via Additive Perturbations (DAAP). Unlike DAMP, DAAP perturbs weights during training with random additive Gaussian noises centered at 00, as detailed in Algorithm 3 in the Appendix. Fig. 7 in the Appendix presents the results of DAMP and DAAP under different noise standard deviations, alongside standard SGD. Overall, Fig. 7 shows that across different experimental settings, the corruption robustness of DAAP is only slightly better than SGD and is worse than DAMP. Therefore, we conclude that MWPs are better than their additive counterparts at improving robustness to corruptions.
@@ 4.3 Benchmark results
In this section, we compare DAMP with Dropout (Srivastava et al., 2014), SAM (Foret et al., 2021), and ASAM (Kwon et al., 2021). For SAM and ASAM, we optimize the neighborhood size $\rho$ by using $10\%$ of the training set as a validation set. Similarly, we adjust the noise standard deviation $\sigma$ for DAMP and the drop rate $p$ for Dropout following the same procedure. For hyperparameters and additional training details, please refer to Appendix F.
@@
Fig. 4 visualizes the predictive errors of DAMP and the baseline methods on CIFAR-10/100 and TinyImageNet, with all methods trained for the same number of epochs. It demonstrates that DAMP consistently outperforms Dropout across various datasets and corruption severities, despite having the same training cost. Notably, DAMP outperforms SAM under most corruption scenarios, even though SAM takes twice as long to train and has higher accuracy on clean images. Additionally, DAMP improves accuracy on clean images over Dropout on CIFAR-100 and TinyImageNet. Finally, ASAM consistently surpasses other methods on both clean and corrupted images, as it employs adversarial MWPs (Section 3). However, like SAM, each ASAM experiment takes twice as long as DAMP given the same epoch counts.
Table 1: DAMP surpasses the baselines on corrupted images in most cases and on average. We report the predictive errors (lower is better) averaged over 3 seeds for the ResNet50 / ImageNet experiments. We evaluate the models on IN-C, IN-$\overline{\mathrm{C}}$, IN-A, IN-D, IN-Sketch, IN-Drawing, IN-Cartoon and adversarial examples generated by FGSM. For FGSM, we use $\epsilon=2/224$. For IN-C and IN-$\overline{\mathrm{C}}$, we report the results averaged over all corruption types and severity levels. The Avg column displays the average of all previous columns except IN-Clean. We use 90 epochs and the basic Inception-style preprocessing for all experiments.
Method Error (%) $\downarrow$
IN-Clean FGSM IN-A IN-C IN-$\overline{\mathrm{C}}$ IN-Cartoon IN-D IN-Drawing IN-Sketch Avg
Dropout $23.6_{\pm 0.2}$ $90.7_{\pm 0.2}$ $\mathbf{95.7_{\pm<0.1}}$ $61.7_{\pm 0.2}$ $61.6_{\pm<0.1}$ $49.6_{\pm 0.2}$ $88.9_{\pm<0.1}$ $77.4_{\pm 1.3}$ $78.3_{\pm 0.3}$ 75.5
DAMP $23.8_{\pm<0.1}$ $\mathbf{88.3_{\pm 0.1}}$ $96.2_{\pm<0.1}$ $\mathbf{58.6_{\pm 0.1}}$ $\mathbf{58.7_{\pm<0.1}}$ $\mathbf{44.4_{\pm<0.1}}$ $88.7_{\pm<0.1}$ $\mathbf{71.1_{\pm 0.5}}$ $\mathbf{76.3_{\pm 0.2}}$ 72.8
SAM $23.2_{\pm<0.1}$ $90.4_{\pm 0.2}$ $96.6_{\pm 0.1}$ $60.2_{\pm 0.2}$ $60.7_{\pm 0.1}$ $47.6_{\pm 0.1}$ $\mathbf{88.3_{\pm 0.1}}$ $74.8_{\pm<0.1}$ $77.5_{\pm 0.1}$ 74.5
ASAM $\mathbf{22.8_{\pm 0.1}}$ $89.7_{\pm 0.2}$ $96.8_{\pm 0.1}$ $58.9_{\pm 0.1}$ $59.2_{\pm 0.1}$ $45.5_{\pm<0.1}$ $88.7_{\pm 0.1}$ $72.3_{\pm 0.1}$ $76.4_{\pm 0.2}$ 73.4
@@
Table 1 presents the predictive errors for the ResNet50 / ImageNet setting on a variety of corruption test sets. It shows that DAMP consistently outperforms the baselines in most corruption scenarios and on average, despite having half the training cost of SAM and ASAM.
Table 2: ViT-S16 / ImageNet (IN) with basic Inception-style data augmentations. Due to the high training cost, we report the predictive error (lower is better) of a single run. We evaluate corruption robustness of the models using IN-A, IN-D, IN-Sketch, IN-Drawing, IN-Cartoon and adversarial examples generated by FGSM. For IN-C and IN-$\overline{\mathrm{C}}$, we report the results averaged over all corruption types and severity levels. For FGSM, we use $\epsilon=2/224$. We also report the runtime of each experiment, showing that SAM and ASAM take twice as long to run than DAMP and AdamW given the same number of epochs. The Avg column displays the average of all previous columns except IN-Clean. DAMP produces the most robust model on average.
Method #Epochs Runtime Error (%) $\downarrow$
IN-Clean FGSM IN-A IN-C IN-$\overline{\mathrm{C}}$ IN-Cartoon IN-D IN-Drawing IN-Sketch Avg
Dropout 100 20.6h 28.55 93.47 93.44 65.87 64.52 50.37 91.15 79.62 88.06 78.31
200 41.1h 28.74 90.95 93.33 66.90 64.83 51.23 92.56 81.24 87.99 78.63
DAMP 100 20.7h 25.50 92.76 92.92 57.85 57.02 44.78 88.79 69.92 83.16 73.40
200 41.1h 23.75 84.33 90.56 55.58 55.58 41.06 87.87 68.36 81.82 70.65
SAM 100 41h 23.91 87.61 93.96 55.56 55.93 42.53 88.23 69.53 81.86 71.90
ASAM 100 41.1h 24.01 85.85 92.99 55.13 55.64 40.74 89.03 67.80 81.47 71.08
ViT-S16 / ImageNet / Basic augmentations